solve for 2x-y=10 and x+5y=49
Answer by oscargut(2103)

solve for 2x-y=10 and x+5y=49
Solution:
2x-y = 10 (1)
x+5y=49 (2)
Doing (1)-2(2)
-11y = -88
y = 8
2x-8=10
2x = 18
x =9
Answer: x=9, y=8